OpenStudy (anonymous):

Niagara Falls in New York is 167 feet high. How long does it take for water to fall from the top to the bottom of Niagara Falls? PLEASE HELP i'm in a chapter involving quadratics

OpenStudy (anonymous):

\[h(t)=-16t^2+167\] when it hits the water, \(h=0\) so solve \[0=167-16t^2\] \[16t^2=167\] \[t^2=\frac{167}{16}\] then take the square root

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

yw